OpenStudy (anonymous):

Express as one logarithm \[3ln(\frac{y^2}{x})-2ln(y)+\frac{1}{3}ln(\frac{x^6y^3}{2})\]

OpenStudy (anonymous):

I got\[ \ln(\frac{y^2\sqrt[3]{2}}{x^5})\]

OpenStudy (anonymous):

\[its \ln(\frac{y^{5}}{x (2^{1/3})})\]

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

gimme me minute!!

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

\[\ln(\frac{y^{6}}{x^{3}})-\ln (y^{2})+\ln(\frac{x^{2}y}{2^1/3})\]

OpenStudy (anonymous):

\[6\ln(\frac{y^{6}*x^{2}*y}{x^{3}*y^{2}*2^{1/3}})\]

OpenStudy (anonymous):

please ignore the 6..

OpenStudy (anonymous):

finally u arrive at the given answer...

OpenStudy (anonymous):

your answer or mine?

OpenStudy (anonymous):

mine....

OpenStudy (anonymous):

cancel the terms....

OpenStudy (anonymous):

But...ln(y^6/x^3)=ln(y^6x^-3)...

OpenStudy (anonymous):

yeah so??

myininaya (myininaya):

\[3\ln(\frac{y^2}{x})-2\ln(y)+\frac{1}{3}\ln(\frac{x^6y^3}{2}) \] Recall \[r \ln(u)=\ln(u^r)\] So first I will do: \[\ln((\frac{y^2}{x})^3)-\ln(y^2)+\ln((\frac{x^6y^3}{2})^\frac{1}{3})\] Now using law of exponents we have \[\ln(\frac{y^6}{x^3})-\ln(y^2)+\ln(\frac{x^2y}{2^\frac{1}{3}})\] So Also recall ln(a)-ln(b)=ln(ab) Giving us: \[\ln(\frac{y^6}{x^3} \div y^2)+\ln(\frac{x^2y}{\sqrt[3]{2}})\] Now recall that a/b divided by c/d =a/b * d/c \[\ln(\frac{y^6}{x^3} \cdot \frac{1}{y^2})+\ln(\frac{x^2 y}{\sqrt[3]{2}})\] Then blah blah I don't now if you guys have reached a result but if you haven't It should be pretty easy to continue from here

myininaya (myininaya):

oops type-o

myininaya (myininaya):

ln(a)-ln(b)=ln(a/b)

OpenStudy (anonymous):

Is that a 3 in the denominator of your first ln?

myininaya (myininaya):

Are you talking the cube or the cube root?

OpenStudy (anonymous):

The answer I got is all the way to the top..

myininaya (myininaya):

Let me know what you don't understand

OpenStudy (anonymous):

why my answer is wrong...

myininaya (myininaya):

I don't know where you went wrong unless you show me your work

OpenStudy (anonymous):

This is what I did \[\ln(\frac{y^6}{x^3})-\ln(y^2)+\ln(\frac{x^2y}{\sqrt[3]{2}})\]\[\ln(y^6x^{-3})-\ln(\frac{x^2y^3}{\sqrt[3]{2}})\]\[\ln(y^6x^{-3})-\ln(\frac{x^2y^3}{\sqrt[3]{2}})\]\[\ln(\frac{y^6x^{-3}\sqrt[3]{2}}{x^2y^3})\]\[\ln(\frac{y^2\sqrt[3]{2}}{x^5})\]

myininaya (myininaya):

So but you are saying 6-3=2 ? let me see if i see any other mistakes

OpenStudy (anonymous):

Okay..

myininaya (myininaya):

and how did you get a minus in front of the last natural log on your 2nd line?

OpenStudy (anonymous):

It was always there, I brought the two second terms together using the product law of logarithms

myininaya (myininaya):

in how did you get a 3 as exponent for that one value of y?

myininaya (myininaya):

ok no no. you did something wrong there then...let me show ya

myininaya (myininaya):

\[\ln(\frac{y^6}{x^3})-\ln(y^2)+\ln(\frac{x^2y}{2^\frac{1}{3}}) \] We agreed to this point right @purplec16 ?

OpenStudy (anonymous):

Yes...

myininaya (myininaya):

I can write it like this this because a+b=b+a addition is commutative \[\ln(\frac{y^6}{x^3})+\ln(\frac{x^2y}{\sqrt[3]{2}})-\ln(y^2)\]

OpenStudy (anonymous):

yup

myininaya (myininaya):

Now you worked with the second two terms first right?

myininaya (myininaya):

So that is what I'm gonna do in this next step

OpenStudy (anonymous):

Yes

myininaya (myininaya):

\[\ln(\frac{y^6}{x^3})+\ln(\frac{x^2 y}{\sqrt[3]{2}} \div y^2)\] by that one property of log : \[\ln(a \div b) \text{ or } \ln(\frac{a}{b})=\ln(a)-\ln(b)\]

myininaya (myininaya):

\[\ln(\frac{y^6}{x^3})+\ln(\frac{x^2 y}{\sqrt[3]{2}} \cdot \frac{1}{y^2})\]

myininaya (myininaya):

do you see what to do from here?

OpenStudy (anonymous):

I guess... am so tired.. I will try to figure it out in the mornign because I am not seeing it but thanks for all your help!

myininaya (myininaya):

which one of the steps I did didn't make sense?

OpenStudy (anonymous):

the 2nd term of your last equation

myininaya (myininaya):

do you know that a/b divided by c/d=a/b * d/c ?

OpenStudy (anonymous):

yes